OpenStudy (anonymous):

What is the equation of the line that contains the point (–5, –1) and has a slope of 4? Write in slope-intercept form. A. y = 4x – 19 B. y = 4x + 19 C. y = 4x – 1 D. y = 4x + 1

OpenStudy (igreen):

First, we can plug it into point-slope form. \(\sf y - y_1 = m(x - x_1)\) Where y1 is the y-value of the point, x1 is the x-value of the point, and 'm' is the slope.

OpenStudy (igreen):

So we have: (-5, -1) Slope of 4 So: (-5, -1) x1 y1 m = 4

OpenStudy (igreen):

Can you plug those values into the equation I gave?

OpenStudy (igreen):

It isn't.

OpenStudy (anonymous):

y=4x-1

OpenStudy (anonymous):

or would it be y=4x+1

OpenStudy (anonymous):

@butterflydreamer

OpenStudy (butterflydreamer):

let's check. So we know for both your questions you have chosen, the slope (m) = 4 so that's correct. Now check by substituting in your point (-5 , -1) So for the first one you chose: " y = 4x - 1" Sub in x = -5 and y = -1 and we get: -1 = 4 ( -5) - 1 simplify -1 = -20 -1 -1 = -21 does -1 equal to -21? I don't think so .. so this is NOT correct :) Next you chose " y = 4x + 1" Sub in x = -5 and y = -1 -1 = 4 ( -5) + 1 -1 = 4 ( -5) + 1 -1 = -20 + 1 -1 = -19 This is also incorrect... SOOOOO we've eliminated two options

OpenStudy (igreen):

@moomoo222 We get this if we plug the numbers in correctly: \(\sf y + 1 = 4(x + 5)\) Can you distribute 4 then subtract 1 to both sides? That will give you your answer.

OpenStudy (texaschic101):

you could just sub your slope(m) and your points (x,y) into y = mx + b...and solve for b...and there is your equation

OpenStudy (anonymous):

so it would be A?

OpenStudy (texaschic101):

or you can do what igreen says and use y - y1 = m(x - x1)

OpenStudy (anonymous):

no it would be c not a

OpenStudy (butterflydreamer):

you can check if A is correct. Sub in x = -5 , y = -1 and m = 4 into that option So option A is : y = 4x – 19 -1 = 4 ( -5) - 19 -1 = -20 - 19 -1 = -39 So not quite right..

OpenStudy (anonymous):

so it is either c or d

OpenStudy (butterflydreamer):

for this question just use the point-slope formula first: y- y1 = m ( x - x1) m = 4 , x1 = -5 , y1 = -1 so we get, y + 1 = 4 ( x +5 ) Then see the "x + 5" in the brackets? We want to multiply both the x and the 5 by 4 So y +1 = 4x + 20 Then just subtract 1 from both sides. (by doing this, we get it into slope - int form : y = mx+b) y + 1 - 1 = 4x + 20 - 1 y = 4x + 19

OpenStudy (anonymous):

that is A

OpenStudy (anonymous):

But a can't be the answer

OpenStudy (anonymous):

it is B

OpenStudy (butterflydreamer):

correct. it's B